Greenhab Report – December 23rd

Greenhab Report
David Murray
23 Dec, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient
Ambient with door open from 9:56 to 15:37
Working Hour: 14:54
Inside temp at working hour: 19 C
Outside temp during working hours: 4.5 C
Inside temperature high: 25 C
Inside temperature low: 15 C
Inside humidity: 16%
Heating
Functioning nominally,
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
N/A
Changes to Crops
Kale and cabbage are sprouting
Morning Research Observations
N/A

Daily Water Usage for Crops
2 gallons
Time(s) of Watering Crops
15:37
Aquaponics
N/A
Narrative
Today I marked the controls and applied the Moringa oleifera fertilizer to 13 tomato plants. The kale and cabbage are starting to sprout but no sign of the spinach. Space in the greenhab is limited so I have to remove some of the tomatoes.
Support/Supplies Needed
N/A

Operations Report – December 23rd

Crew 185

Operations Report Saturday 12/23/2017

SOL: 5

Name of person filing report: A.Passeron

Non-nominal systems:

– Generator :
The generator had a failure during the night, Fault message reading "Low battery". Jump-start from Shannon’s Shuttle battery at did not solve the issue. Generator had a second failure at rougthly 12:00pm. Second jump-start done and generator turned on at 5:30pm. The generator is currently running.
At 6:20pm, settings are :
+ Oil pressure : 45 psi
+ Speed : 1815 rpm
+ Battery charge : Fluctuating between 8.7V and 9.0V
At 6:40pm, settings are :
+ Oil pressure : 45 psi
+ Speed : 1815 rpm
+ Battery charge : Fluctuating between 7.8V and 8.0V
The charge of the battery is thus plummeting. The battery of the generator is not recharged while generator is on.

– Water heater tests :
I went to the shower and did the following tests.

@8:42am first trial with nominal settings (right dial on the second top black bar) :
hot water turned on all the way for 2 minutes, initial temperature 51F, final temperature 64F.
@9:33am second trial with max temperature (right dial turned to the top black bar, but I can go further) :
hot water turned on all the way for 4 minutes, initial temperature 55F, final temperature 64F.

– Water shortage :
Crew aware of concern about water shortage. We are now using jerrycans for drinking water. We tried to pump as much water as possible from the trailer tanks into the static tank. We use the shallow water of trailer tanks to water the plants of the GreenLab.

Generator (hours run):
Generator turned off at unknown time (failure during the night)
Jump-start done at 8:10am
Generator turned on at 8:10am
Generator turned off at roughtly 12:00pm (second failure during the EVA)
Jump-start done at 5:30pm
Generator turned on at 5:30pm

Solar— SOC
@ 8:00am : 57%
@ 9:20am : 60%
@ 10:00am : 64%
@ 11:00am : In EVA
@ 12:00pm : In EVA
@ 1:00pm : In EVA
@ 2:00pm : 100%
@ 3:00pm : 98%
@ 4:50pm : 90%
@ 5:00pm : 88%
@ 5:30pm : 84%
@ 6:40pm : 82%

Diesel – 78%

Propane – 52%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2.8 Gallons

Water (trailer) – 20 gallons

Water (static) – 280 gallons

Trailer to Static Pump used – Yes

Static to Loft Pump used – Yes

Water Meter: 50 gallons

Toilet tank emptied: Yes

ATVs Used: Yamaha blue Grizzly 350 #3
Oil Added? Yes
ATV Fuel Used:
# Hours the ATVs were used today: 02:00 hours
Notes on ATVs: ATVs were nominal.

Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes

Spirit rover used: Yes
Hours: 02:00 Hours
Beginning charge: 100%
Ending charge: 60%
Currently charging: Yes

Opportunity rover used: No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:

Curiosity rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

HabCar used and why, where? No

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: Water shortage prevention. Water heater doesn’t seem to work properly.

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support:

EVA Report – December 23rd

Report: Crew 185 EVA #4 for SOL 5.
Propose of EVA: Collecting geological samples for mobility testing. Then, walking with the EVA suits and discover the terrain around our position.

EVA Report:
(i) The first 40 min were dedicated to the engineering check during which we have filled the static tank with the left unfrozen water of the tanks on the trailer. Then, using a pan, we collected other (frozen) water that can be used for the Green Hab.
(ii) Arrived at the White Moon, we did a short walk around. Because strangers were near the vehicles, we went back, and we collected a few geological samples.
(iii) Then, we stopped about 1 km from the Hab to locate recognition signs to be used for the following two EVAs.
(iv) Back at the Hab, we refilled the ATV 350 and plugged the rover.

EVA #5:
Start time: 1030am
End time: 1230pm -> finished at 2pm

Ilaria Cinelli: Leader
John : Astro 1
Arno: Astro 2
Destination: White Moon (110.8W,38.44N).
Vehicles: 1 ATV (Ilaria), 1 Rover (John and Arno)
CapCom(s): David and Thibo

Written by Commander Ilaria Cinelli

Sol Summary – December 23rd

Crew 185 Sol Summary Report 23 Dec 2017

Sol 5

Beginning of a Navigation Exercise

David Murray

Mission Status: We had slight generator problems but everything is under control and we still have charge.

Sol Activity Summary: The crew ate breakfast and disbanded for individual tasks before meeting back up to do the EVA prep. Ilaria, John, and Arno made up the EVA crew today. They placed objects that the next EVA crew will have to locate in the field based on instructions from CapComm. After the EVA Arno and I participated in John’s medical training.

Look Ahead Plan: Tomorrow we’re planning to do a split EVA to test the X-1 and to locate the items that today’s EVA crew put in the field.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: Healthy

EVA: 3 person EVA to use ATV’s and hide objects for a navigation study.

Reports to be filed: Journalist Report, Greenhab Report, EVA Request, EVA Report, Sol Summary

Support Requested: None

Journalist Report – December 22nd

[Sol 4]

And we are back to red! After 24h of snow and landscapes looking more like Europa than Mars, the sun is shinning again and the temperatures are positive.

Today, we did a short walking EVA to pick up geological samples around the Hab and to see if X-1 is capable of doing it. We brought back a dozen interesting rocks and samples of soil.

Then, this afternoon, we practiced CPR with John to be sure that we would be able to do it if needed. We all managed to rescue Phil the mannequin successfully!

The crew is doing great so far. The only concern is that our water supplies are running low, even though we are careful and didn’t have a shower yet. We discussed tonight how to make it even lower. Hopefully, we will have enough water for the rest of the mission. Think of us when you will have your next shower, it is a dream here!

Greetings from the Red Planet,

Thibault, ExO and spacesuit engineer for the Crew 185

Greenhab Report – December 22nd

Greenhab Report
David Murray
22 Dec, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient
Ambient with door open from about noon to 15:30
Working Hour: 15:30
Inside temp at working hour: 24 C
Outside temp during working hours: -0.5 C
Inside temperature high: 33 C
Inside temperature low: 14 C
Inside humidity: 16%
Heating
Functioning nominally,
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
N/A
Changes to Crops
None
Morning Research Observations
N/A

Daily Water Usage for Crops
4 gallons + 3 gallons of miracle grow and water
Time(s) of Watering Crops
15:30
Aquaponics
N/A
Narrative
I transplanted some more tomatoes and applied Miracle Grow fertilizer to all the plants in the greenhab with the exception of my control and experimental groups of tomato plants. I added tomato cages to the bean plants for some support because that’s the best option I have right now. All the plants are looking healthy.
Support/Supplies Needed
N/A

Operations Report – December 22nd

Crew 185

Operations Report Friday 12/22/2017

SOL: 4

Name of person filing report: A.Passeron

Non-nominal systems:

-Spacesuit N°3 :
The ventilation system worked properly during the whole EVA. The running time of the fully charged battery is thus more than 1h20min.

– Shower-head :
Shower-head leak was fixed at 9:10am.

– Water shortage :
Crew aware of concern about water shortage. We are now using jerrycans for drinking water.

Generator (hours run):
Generator turned off at 8:10 am (SOC at 100%)
Generator turned on at 5:00pm (SOC at 90%)

Solar— SOC
@ 8:10am : 100%
@ 9:06am : 95%
@ 10:00am : 100%
@ 11:00am : did not check because of EVA
@ 12:00pm : 100%
@ 1:05pm : 100%
@ 2:35pm : 100%
@ 3:30pm : 99%
@ 5:00pm : 90%
@ 5:40pm : 90%

Diesel – 80%

Propane – 54%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 3.5 Gallons

Water (trailer) – 50 gallons

Water (static) – 280 gallons

Trailer to Static Pump used – No

Static to Loft Pump used – Yes

Water Meter: 50 gallons

Toilet tank emptied: No

ATVs Used: No
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.

Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes

Opportunity rover used: No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes

Curiosity rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

HabCar used and why, where? No

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: Shower-head leak is fixed. Water shortage prevention.

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support:

GreenHab Report – December 21st

Greenhab Report
David Murray
21 Dec, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient
Working Hour: 14:06
Inside temp at working hour: 18 C
Outside temp during working hours: -0.5 C
Inside temperature high: 20 C
Inside temperature low: 14 C
Inside humidity: 16%
Heating
Functioning nominally,
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
N/A
Changes to Crops
None
Morning Research Observations
N/A
Daily Water Usage for Crops
2 gallons
Time(s) of Watering Crops
14:06
Aquaponics
N/A
Narrative
We had snow today so the greenhab stayed relatively cool with a high of 20 C, the same temperature that the heater shuts off at. With a limited water supply, the cool temperatures were welcome as I was able to use 1/3 of the water that I normally do on the plants. No signs of the seeds sprouting but I wouldn’t expect it at this point anyway. Tomorrow I’ll transplant more tomatoes and select plants to use as the control and test subject of the Moringa fertilizer experiment. Because the duration of this experiment continues into the next crew, I’d like to set the time for reapplication around a week after the next crew arrives so they have time to get settled and comfortable.
Support/Supplies Needed
N/A

Operations Report – December 21st

Crew 185
Operations Report Thursday 12/21/2017
SOL: 3
Name of person filing report: A.Passeron
Non-nominal systems:
-Spacesuit N°3 :
The ventilation system stopped once again during the engineering EVA. I opened the backpack at 3:00pm and tried to find a faulty connection with the ohmmeter. All the wires are well connected. The faulty connection may thus be in the outlet itself. Still, the backpack can be charged by slightly moving the plug. The ventilation system is operational. Tomorrow, I will check the running time of the fully-charged spacesuit to test whether or not the battery is defective.
– Outer window of front airlock and central window of living area:
Windows covered with bubblewrap to insulate against cold.
– Water heater :
The water heater is plugged. The front screen displays 061 to 063. We haven’t touched anything except the plug yet.
At 8:20am, I went to the shower and turned the hot water on all the way for 1 minute but water remained cold.
In the living area, water does not flow into the sink when hot is turned on.
– Shower-head :
Shower-head leaks when turned on.
Generator (hours run):
Generator turned off at 8:00 am (SOC at 78%)
Generator turned on at 8:55am because of snowy weather (SOC at 69%)
Solar— SOC @ 5:05pm : 100%
Diesel – 80%
Propane – 56%
Ethanol Free Gasoline (5 Gallon containers for ATV) – 3.5 Gallons
Water (trailer) – 50 gallons
Water (static) – 320 gallons
Trailer to Static Pump used – No
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: Yes
ATVs Used: No
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Opportunity rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Curiosity rover used:  No
Hours:
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: The outlet of spacesuit N°3 has a faulty connection. Checking the battery lifespan tomorrow.
Summary of Hab operations: Front airlock and living area windows are covered with bubblewrap. Shower-head leaks when turned on. Still trying to operate the water heater.
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support: